OpenStudy (anonymous):

A line has a slope of 9/4 and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct. @ganeshie8 @campbell_st @mathstudent55

OpenStudy (danjs):

y=m*x+b y=(9/4)x-14

OpenStudy (anonymous):

A. -9x + 4y = -56 B. 9x + 4y = 56 C. y = 9/4x - 14 D. y = 9/4x + 14

OpenStudy (danjs):

then just play around with that and put it in other forms

OpenStudy (anonymous):

is a one too

OpenStudy (danjs):

y=(9/4)x-14 4y = 9x - 56 -9x +4y = -56

OpenStudy (anonymous):

so its a and c

OpenStudy (danjs):

correct

OpenStudy (anonymous):

can you help me with a couple more

OpenStudy (danjs):

(y2-y1)=m(x2-x1) Point slope form y=mx+b ax+by=0 standard form

OpenStudy (danjs):

yes sure

OpenStudy (anonymous):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (danjs):

the point (1,8) ?

OpenStudy (anonymous):

A line has a slope of 3 and contains the point (−1, −8).

OpenStudy (danjs):

Given: m=3 and p(x,y)=(1,8) y=mx+b 8=(3)(1)+b b= 8/3 so, y= 3*x + 8/3

OpenStudy (danjs):

or 3y = 9x +8 -9x + 3y = 8

OpenStudy (danjs):

not sure if that is correct, your copy paste is messed up

OpenStudy (anonymous):

ok but can you help me with some more

OpenStudy (danjs):

sure

OpenStudy (anonymous):

Cheryl has 16 cupcakes made. She can make 48 cupcakes an hour. A customer asks for 160 cupcakes. How many hours will it take Cheryl to make enough cupcakes for the customer? i think ist 3

OpenStudy (danjs):

i used (1,8) , replace that with (-1,-8)

OpenStudy (danjs):

She already has 16 and can make 48 each hour and needs 160 160 = 16 + 48*(#hours)

OpenStudy (anonymous):

is it 3

OpenStudy (danjs):

yep

OpenStudy (anonymous):

Gary adds $12 to his savings each week. At Week 3, Gary had $60 total in his savings. How much will Gary have saved after 15 wk?

OpenStudy (anonymous):

i think its 204 @DanJS

OpenStudy (danjs):

Total = 12*(#week) + 24

OpenStudy (danjs):

because, at week 3 he had 60, 12*3 = 36 so he started at week 0 (y-intercept) with 24 dollars and it is a linear equation

OpenStudy (danjs):

total = starting amount + 12 * # weeks